OpenStudy (anonymous):

Rational Functions: 4/6-x + x^2/6-x = 2

OpenStudy (anonymous):

Might help to write it symbolically:\[\frac{4}{6-x}+\frac{x^2}{6-x}=2\] \[\frac{4+x^2}{6-x}=2\] \[4+x^2=12-2x\] \[x^2+2x-8=0\] \[(x+4)(x-2)=0\] x = -4 or x = 2. Tell me if any part of that was confusing.

OpenStudy (anonymous):

Nope, I get it I had a feeling that was the answer just wanted to double check. Thank You

OpenStudy (anonymous):

Great, you're welcome.

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

1/x^2-7x+10 = x/x-5+1/x-2

OpenStudy (anonymous):

\[\frac{1}{x^2-7x+10} = \frac{x}{x-5}+\frac{1}{x-2}\] \[\frac{1}{x^2-7x+10} = \frac{x^2-x-5}{x^2-7x+10}\] \[0= \frac{x^2-x-6}{x^2-7x+10}\] \[x^2-x-6 = 0\] x = 3 or -2. But you have to plug back in to see if there's any problems - turns out, there aren't. Those are the solutions.

OpenStudy (anonymous):

Sweet! Thank You again. Happy Holidays

OpenStudy (anonymous):

Happy Holidays!